# A new HOC scheme for computation of flow and heat transfer in nonuniform polar grids

### v2

@@ 2 Mathematical formulation and discretization procedure
$B_{2}=$ $\frac{1}{(1+\alpha_{r})h_{r_{i-1}}^{2}}\left(\frac{4A_{1}}{\alpha_{r}}+A_{3}\right),\qquad\qquad\qquad B_{3}=\frac{1}{(1+\alpha_{r})h_{r_{i-1}}^{2}}\left(4A_{1}-A_{3}\right),$
$B_{4}=$ $\frac{1}{(1+\alpha_{\theta})h_{\theta_{j-1}}^{2}r_{i}^{2}}\left(\frac{4A_{2}}{\alpha_{\theta}}+A_{4}\right),\qquad\qquad\quad B_{5}=\frac{1}{(1+\alpha_{\theta})h_{\theta_{j-1}}^{2}r_{i}^{2}}\left(4A_{2}-A_{4}\right),$
$B_{6}=$ $\frac{A_{1}}{(1+\alpha_{r})h_{r_{i-1}}},\qquad\qquad\qquad\qquad\qquad\qquad B_{7}=\frac{A_{3}}{h_{r_{i-1}}},$
$B_{8}=$ $\frac{A_{2}}{(1+\alpha_{\theta})h_{\theta_{j-1}}r_{i}^{2}},\qquad\qquad\qquad\qquad\qquad\quad B_{9}=\frac{A_{3}}{h_{\theta_{j-1}}r_{i}^{2}}.$
Here, the coefficients $B_{i}$s, $i\in{1,2,3,\dots,9}$ of the algebraic expression depends entirely on the values of grid spacing and they remain unchanged throughout the computation once the grid is set up. Therefore, for situations with constant convection and diffusion coefficients the current scheme also offers the inherent benefit of dealing with a resulting system of equations containing constant coefficients only. Further, one can easily notice that the discrete operator $\mathcal{A}$ also contains the radial and tangential derivatives of the flow variables, which need to be approximated up to the appropriate order of accuracy. In this aspect, we generalize the Padé scheme [12] to obtain the following approximations of the spatial derivatives on nonuniform polar grids.
$\left(1+\frac{h_{r_{i}}h_{r_{i-1}}}{6}\delta_{rr}\right)\phi_{r_{i,j}}=$ $\left(\delta_{r}-\frac{h_{r_{i}}-h_{r_{i-1}}}{2}\delta_{rr}\right)\phi_{i,j}$ (12)
$+\mathcal{O}\left((h_{r_{i}}-h_{r_{i-1}})(h_{r_{i}}^{2}-h_{r_{i}}h_{r_{i-1}}+h_{r_{i-1}}^{2})\right).$
@@
$\left(1+\frac{h_{\theta_{j}}h_{\theta_{j-1}}}{6}\delta_{\theta\theta}\right)\phi_{\theta_{i,j}}=$ $\left(\delta_{\theta}-\frac{h_{\theta_{j}}-h_{\theta_{j-1}}}{2}\delta_{\theta\theta}\right)\phi_{i,j}$ (13)
$+\mathcal{O}\left((h_{\theta_{j}}-h_{\theta_{j-1}})(h_{\theta_{j}}^{2}-h_{\theta_{j}}h_{\theta_{j-1}}+h_{\theta_{j-1}}^{2})\right).$
Subsequently, we obtain the corresponding algebraic system for equations 12 and 13 as,
@@
$\phi_{r_{1,j}}=$ $-\frac{h_{r_{1}}}{h_{r_{2}}(h_{r_{1}}+h_{r_{2}})}\phi_{3,j}+\frac{h_{r_{1}}+h_{r_{2}}}{h_{r_{1}}h_{r_{2}}}\phi_{2,j}-\frac{2h_{r_{1}}+h_{r_{2}}}{h_{r_{1}}(h_{r_{1}}+h_{r_{2}})}\phi_{1,j}$
$+\mathcal{O}\left(h_{r_{2}}(h_{r_{1}}+h_{r_{2}})^{2}\right),$ (16a)
$\phi_{r_{n_{r},j}}=$ $\frac{h_{r_{n_{r}-1}}}{h_{r_{n_{r}-2}}(h_{r_{n_{r}-1}}+h_{r_{n_{r}-2}})}\phi_{n_{r}-2,j}-\frac{h_{r_{n_{r}-1}}+h_{r_{n_{r}-2}}}{h_{r_{n_{r}-1}}h_{r_{n_{r}-2}}}\phi_{n_{r}-1,j}$
$+\frac{2h_{r_{n_{r}-1}}+h_{r_{n_{r}-2}}}{h_{r_{n_{r}-1}}(h_{r_{n_{r}-1}}+h_{r_{n_{r}-2}})}\phi_{n_{r},j}+\mathcal{O}\left(h_{r_{n_{r}-1}}(h_{r_{n_{r}-1}}+h_{r_{n_{r}-2}})^{2}\right),$ (16b)
@@
$\phi_{\theta_{i,1}}=$ $-\frac{h_{\theta_{1}}}{h_{\theta_{2}}(h_{\theta_{1}}+h_{\theta_{2}})}\phi_{i,3}+\frac{h_{\theta_{1}}+h_{\theta_{2}}}{h_{\theta_{1}}h_{\theta_{2}}}\phi_{i,2}-\frac{2h_{\theta_{1}}+h_{\theta_{2}}}{h_{\theta_{1}}(h_{\theta_{1}}+h_{\theta_{2}})}\phi_{i,1}$
$+\mathcal{O}\left(h_{\theta_{2}}(h_{\theta_{1}}+h_{\theta_{2}})^{2}\right),$ (17a)
$\phi_{\theta_{i,n_{\theta}}}=$ $\frac{h_{\theta_{n_{\theta}-1}}}{h_{\theta_{n_{\theta}-2}}(h_{\theta_{n_{\theta}-1}}+h_{\theta_{n_{\theta}-2}})}\phi_{i,n_{\theta}-2}-\frac{h_{\theta_{n_{\theta}-1}}+h_{\theta_{n_{\theta}-2}}}{h_{\theta_{n_{\theta}-1}}h_{\theta_{n_{\theta}-2}}}\phi_{i,n_{\theta}-1}$
$+\frac{2h_{\theta_{n_{\theta}-1}}+h_{\theta_{n_{\theta}-2}}}{h_{\theta_{n_{\theta}-1}}(h_{\theta_{n_{\theta}-1}}+h_{\theta_{n_{\theta}-2}})}\phi_{i,n_{\theta}}+\mathcal{O}\left(h_{\theta_{n_{\theta}-1}}(h_{\theta_{n_{\theta}-1}}+h_{\theta_{n_{\theta}-2}})^{2}\right).$ (17b)
@@ 3 Solution of algebraic system
Similarly matrix representation of equation 14 and equation 15 are
@@ 4 Numerical examples
To examine the accuracy and effectiveness of the present scheme on nonuniform polar grids, it has been applied to six different problems of varying complexities. Validation study is carried out by solving the convection-diffusion of the Gaussian pulse. We study the efficiency of the scheme in handling complex flow patterns by solving the driven polar cavity problem. Subsequently, we enter the field of heat transfer by tackling the problem of heat convection in an annulus. Finally, we carry out a comprehensive study of flow and heat transfer around a circular cylinder. The intention behind selecting these problems is to highlight the inherent flexibility of nonuniform grids whereby certain portions of the solution domain can be better resolved with grid clustering. Furthermore, the overwhelming number of numerical solutions present in the literature gives us the leverage to compare our numerical solution with the existing ones establishing the robustness of this newly developed scheme. All the computations are executed on an Intel i7-based PC with 3.40 GHz CPU and 32 GB RAM. The tolerance for inner and outer iterations is set to be 1.0e-10.
@@ 4.2 Problem 2: Driven polar cavity
A driven flow inside a polar cavity is also been simulated using the present scheme. This problem is governed by the transient form of N-S equation which can be given as,
@@
The problem is defined in an annular region $\left[1,2\right]\times\left[\dfrac{\pi-1}{2},\dfrac{\pi+1}{2}\right]$. A schematic illustration of the problem is presented in fig. 3. The region $ABCD$ is the domain of the problem: the height $AB=DC$ of the cavity is equal to the radius $R_{0}$=1 of the inner circle. The inner circular arc $DA$ rotates with a unit velocity in the clockwise direction which drives the flow inside the cavity. All the other boundary walls are stationary throughout the simulation. From this, we can derive the boundary conditions for $u$ and $v$ as: $u=0$, $v=-1$ at the moving boundary $DA$ and $u$=0, $v$=0 at all the other boundaries. The streamfunction values are set to be zero at all the boundaries $(\psi=0)$. For this problem, the Reynolds number is defined as $Re=\dfrac{v_{0}R_{0}}{\nu}$, where $v_{0}$ is magnitude of the surface velocity of rotating side $DA$ and $\nu$ is the kinematic viscosity of the fluid. This is imperative to mention here that all boundary walls follow no-slip condition with the fluid inside the cavity, which supplies the following Neumann boundary condition for vorticity.
@@
The computations for this problem are carried out for Reynolds numbers 55, 350, 1000, 2000, 3000 and 5000. The post processed steady-state streamfunction $(\psi)$ and vorticity $(\omega)$ contours for $55\leq Re\leq 3000$ are plotted in fig. 4. As $Re$ increases the primary vortex is seen to move gradually towards the right wall of the cavity. Similar to the case of flow inside a square cavity, with an increase in the value of $Re$ the secondary vortices at the corners opposite to the moving wall grow in size, particularly the secondary left vortex gains significantly more size as compared to the secondary right vortex and for higher values of $Re$ the former tends to occupy the entire left side of the cavity. Further, at $Re$=3000 a tertiary vortex is seen to form at the top left corner of the cavity. It is to be mentioned that all the findings are in accordance with those available in the literature [16, 9, 10, 17, 11] both experimentally and numerically. Additionally, at higher $Re$ values the vorticity contours move away from the center of the cavity toward the cavity walls thereby developing very strong vorticity gradients in the vicinity of the cavity walls.
@@ 55 $33^{2}$
Present -3.559e-3 0.714 1.770 -7.062e-3 -0.452 1.149
a
b
Figure 10: Problem 2: (a) Time history and (b) power spectrum of total energy of the flow inside a polar cavity for $Re=5000$.
Subsequently, we carry out our computation for $Re=5000$ on a $129\times 129$ nonuniform polar grid. To the best of the authors knowledge, the solution to this problem for Reynolds numbers as high as 5000 has not been computed using a transient scheme. Sen and Kalita [17] in their work used a steady code to arrive at a converged solution at $Re=5000$. Also authors used a $82\times 118$ grid with exponential clustering at the inner wall along the radial direction. In this work we intend to better capture the physics of the flow and hence use appropriately clustering on all boundaries. It is interesting to notice that the present discretization of the N-S equation shows a gradual convergence towards a stable periodic solution rather than a steady-state solution as reported in [17]. Time evolution of the physical property of the flow inside the cavity helps to establish the periodic nature of the solution. Temporal progression of the total energy inside the cavity is presented in fig. 10. We perform the spectral density analysis for the variation of energy which results in a single frequency peak as can be seen in fig. 10. This further ascertains the periodicity of the flow. Nevertheless, accurate estimation of the frequency and time-period of the solution leaves further scope of investigation of this problem.
@@ 4.3 Problem 3: Natural convection in horizontal concentric annulus
We further analyze the efficiency of the newly developed scheme to tackle heat transfer by implementing it to solve natural heat convection inside a horizontal concentric annulus. This problem has gained a considerable amount of attention from researchers over the years because of its relevance and applicability in various engineering and physical situations, such as heat exchangers, solar collectors, nuclear reactors, thermal energy storage systems, etc [19, 20, 21, 22, 23, 24, 25, 26, 27]. It is governed by the transient Boussinesq equations which in the polar coordinate system are given as,
@@
The problem setup consists of two concentric circular walls of which the inner circle has a radius $R_{i}$ and the outer circle has a radius $R_{o}$. Two different temperatures are maintained at the inner and outer circular walls of the annulus as shown in fig. 11. The simulation is carried out for $Ra$ values 2.38e3, 9.50e3, 4.70e4, 6.19e4, and 1.02e5. The values of the $Pr$ is set as 0.706, 0.717, and 0.718. For each case, we have considered $L/D_{i}$=0.8, where $L=R_{o}-R_{i}$ and $D_{i}=2\texttimes R_{i}$. The inner wall is heated to a unit nondimensional temperature $(\mathcal{T}_{h}=1)$ keeping the outer wall’s temperature at zero $(\mathcal{T}_{c}=0)$. This temperature difference serves as the driving force for the flow in this simulation. The values of streamfunction and vorticity at the inner and outer boundaries of the annulus are obtained respectively from no-slip criteria between the fluid and the walls of the annulus and one-sided approximations introduced in equations 28, 29, 30 and 31. The prior studies help us to identify that the near boundary region of the inner circle and the radial line $\theta=\dfrac{\pi}{2}$ should be observed cautiously. This motivates us to work with a mesh where grid points are clustered in these regions (see fig. 11). The grid is generated by making adequate changes in equations 32 and 22 and choosing the associated parameters as $L_{r}=0.8$, $L_{\theta}=2\pi$, $\Theta_{r}=\pi$, $\Theta_{\theta}=2\pi$, $\lambda_{r}=-0.4$ and $\lambda_{\theta}=0.25$.
@@
The numerical results of the present simulation have been presented in figs. 12 and 13. We carry out a qualitative comparison in fig. 12 where the numerically obtained steady-state isotherm contours for $Ra=4.7e4$, $Pr$=0.706 are depicted alongside that of the experimental study [21]. The figure shows that the fluid near the hot inner wall of the annulus flows up along the inner boundary and a strong plume emerges above the inner boundary at $\theta=\dfrac{\pi}{2}$, while the fluid near the cold outer wall flows down to the bottom of the annuli as an effect of buoyancy caused by the temperature difference. Once the flow reaches a steady state, the temperature field becomes symmetric. The maximum temperature gradient appears in the raised plume. Good agreements are obtained between the experimental and numerical solutions. Fig. 13 contains the steady-state isotherms alongside the streamfunctions for different $Ra$ values. The figure shows that for lower $Ra$ values natural convection is dominated by heat conduction. Almost concentric isotherms can be seen for $Ra=2.38e3$ while two symmetric vortices can be seen caused by the weak natural convection. As the value of $Ra$ increases the fluid motion driven by buoyancy force also increases, leading to a stronger convection. As a result, the isotherms start moving upward and change shape to form a plume. The plume becomes more prominent with a further increase in the $Ra$ value.

### v3

@@ 2 Mathematical formulation and discretization procedure
$B_{2}=$ $\frac{1}{(1+\alpha_{r})h_{r_{i-1}}^{2}}\left(\frac{4A_{1}}{\alpha_{r}}+A_{3}\right),\qquad\quad\quad B_{3}=\frac{1}{(1+\alpha_{r})h_{r_{i-1}}^{2}}\left(4A_{1}-A_{3}\right),$
$B_{4}=$ $\frac{1}{(1+\alpha_{\theta})h_{\theta_{j-1}}^{2}r_{i}^{2}}\left(\frac{4A_{2}}{\alpha_{\theta}}+A_{4}\right),\quad\quad\quad B_{5}=\frac{1}{(1+\alpha_{\theta})h_{\theta_{j-1}}^{2}r_{i}^{2}}\left(4A_{2}-A_{4}\right),$
$B_{6}=$ $\frac{A_{5}}{(1+\alpha_{r})h_{r_{i-1}}},\qquad\qquad\qquad\qquad\qquad B_{7}=\frac{A_{3}}{h_{r_{i-1}}},$
$B_{8}=$ $\frac{A_{6}}{(1+\alpha_{\theta})h_{\theta_{j-1}}r_{i}^{2}},\qquad\qquad\qquad\qquad\quad B_{9}=\frac{A_{4}}{h_{\theta_{j-1}}r_{i}^{2}}.$
Here, the coefficients $B_{i}$’s, $i\in\{1,2,3,\dots,9\}$ of the algebraic expression depends entirely on the values of grid spacing and they remain unchanged throughout the computation once the grid is set up. Therefore, for situations with constant convection and diffusion coefficients the current scheme also offers the inherent benefit of dealing with a resulting system of equations containing constant coefficients only. Further, one can easily notice that the discrete operator $\mathcal{A}$ also contains the radial and tangential derivatives of the flow variables, which need to be approximated up to the appropriate order of accuracy. In this aspect, we generalize the Padé scheme [12] to obtain the following approximations of the spatial derivatives on nonuniform polar grids.
$\left(1+\frac{h_{r_{i}}h_{r_{i-1}}}{6}\delta_{rr}\right)\phi_{r_{i,j}}=\left(\delta_{r}-\frac{h_{r_{i}}-h_{r_{i-1}}}{2}\delta_{rr}\right)\phi_{i,j}+\mathcal{O}\left((h_{r_{i}}-h_{r_{i-1}})(3h_{r_{i}}^{2}-h_{r_{i}}h_{r_{i-1}}+3h_{r_{i-1}}^{2})\right).$ (15)
@@
$\left(1+\frac{h_{\theta_{j}}h_{\theta_{j-1}}}{6}\delta_{\theta\theta}\right)\phi_{\theta_{i,j}}=\left(\delta_{\theta}-\frac{h_{\theta_{j}}-h_{\theta_{j-1}}}{2}\delta_{\theta\theta}\right)\phi_{i,j}+\mathcal{O}\left((h_{\theta_{j}}-h_{\theta_{j-1}})(3h_{\theta_{j}}^{2}-h_{\theta_{j}}h_{\theta_{j-1}}+3h_{\theta_{j-1}}^{2})\right).$ (16)
Subsequently, we obtain the corresponding algebraic system for equations 15 and 16 as,
@@
$\phi_{r_{1,j}}=$ $-\frac{h_{r_{1}}}{h_{r_{2}}(h_{r_{1}}+h_{r_{2}})}\phi_{3,j}+\frac{h_{r_{1}}+h_{r_{2}}}{h_{r_{1}}h_{r_{2}}}\phi_{2,j}-\frac{2h_{r_{1}}+h_{r_{2}}}{h_{r_{1}}(h_{r_{1}}+h_{r_{2}})}\phi_{1,j}+\mathcal{O}\left(h_{r_{2}}(h_{r_{1}}+h_{r_{2}})\right),$ (19a)
$\phi_{r_{n_{r},j}}=$ $\frac{h_{r_{n_{r}-1}}}{h_{r_{n_{r}-2}}(h_{r_{n_{r}-1}}+h_{r_{n_{r}-2}})}\phi_{n_{r}-2,j}-\frac{h_{r_{n_{r}-1}}+h_{r_{n_{r}-2}}}{h_{r_{n_{r}-1}}h_{r_{n_{r}-2}}}\phi_{n_{r}-1,j}+\frac{2h_{r_{n_{r}-1}}+h_{r_{n_{r}-2}}}{h_{r_{n_{r}-1}}(h_{r_{n_{r}-1}}+h_{r_{n_{r}-2}})}\phi_{n_{r},j}$
$+\mathcal{O}\left(h_{r_{n_{r}-1}}(h_{r_{n_{r}-1}}+h_{r_{n_{r}-2}})\right),$ (19b)
@@
$\phi_{\theta_{i,1}}=$ $-\frac{h_{\theta_{1}}}{h_{\theta_{2}}(h_{\theta_{1}}+h_{\theta_{2}})}\phi_{i,3}+\frac{h_{\theta_{1}}+h_{\theta_{2}}}{h_{\theta_{1}}h_{\theta_{2}}}\phi_{i,2}-\frac{2h_{\theta_{1}}+h_{\theta_{2}}}{h_{\theta_{1}}(h_{\theta_{1}}+h_{\theta_{2}})}\phi_{i,1}+\mathcal{O}\left(h_{\theta_{2}}(h_{\theta_{1}}+h_{\theta_{2}})\right),$ (20a)
$\phi_{\theta_{i,n_{\theta}}}=$ $\frac{h_{\theta_{n_{\theta}-1}}}{h_{\theta_{n_{\theta}-2}}(h_{\theta_{n_{\theta}-1}}+h_{\theta_{n_{\theta}-2}})}\phi_{i,n_{\theta}-2}-\frac{h_{\theta_{n_{\theta}-1}}+h_{\theta_{n_{\theta}-2}}}{h_{\theta_{n_{\theta}-1}}h_{\theta_{n_{\theta}-2}}}\phi_{i,n_{\theta}-1}+\frac{2h_{\theta_{n_{\theta}-1}}+h_{\theta_{n_{\theta}-2}}}{h_{\theta_{n_{\theta}-1}}(h_{\theta_{n_{\theta}-1}}+h_{\theta_{n_{\theta}-2}})}\phi_{i,n_{\theta}}$
$+\mathcal{O}\left(h_{\theta_{n_{\theta}-1}}(h_{\theta_{n_{\theta}-1}}+h_{\theta_{n_{\theta}-2}})\right).$ (20b)
@@ 3 Solution of algebraic system
Similarly, matrix representation of equation 17 and equation 18 are
@@ 4 Numerical examples
To examine the accuracy and effectiveness of the present scheme on nonuniform polar grids, it has been applied to four different problems of varying complexities. Validation study is carried out by solving the convection-diffusion of the Gaussian pulse. We study the efficiency of the scheme in handling complex flow patterns by solving the driven polar cavity problem. Subsequently, we enter the field of heat transfer by tackling the problem of heat convection in an annulus. Finally, we carry out a comprehensive study of flow and heat transfer around a circular cylinder. The intention behind selecting these problems is to highlight the inherent flexibility of nonuniform grids whereby certain portions of the solution domain can be better resolved with grid clustering. Furthermore, the overwhelming number of numerical solutions present in the literature gives us the leverage to compare our numerical solution with the existing ones establishing the robustness of this newly developed scheme. All the computations are executed on an Intel i7-based PC with 3.40 GHz CPU and 32 GB RAM. The tolerance for inner and outer iterations is set to be $101.0e-$ 10.
@@ 4.2 Problem 2: Driven polar cavity
A driven flow inside a polar cavity is also been simulated using the present scheme. This problem is governed by the transient form of N-S equations which can be given as,
@@
The problem is defined in an annular region $\left[1,2\right]\times\left[\dfrac{\pi-1}{2},\dfrac{\pi+1}{2}\right]$. A schematic illustration of the problem is presented in fig. 3. The region $ABCD$ is the domain of the problem: the height $AB=DC$ of the cavity is equal to the radius $R_{0}$=1 of the inner circle. The inner circular arc $DA$ rotates with a unit velocity in the clockwise direction which drives the flow inside the cavity. All the other boundary walls are stationary throughout the simulation. From this, we can derive the boundary conditions for $u$ and $v$ as: $u=0$, $v=-1$ at the moving boundary $DA$ and $u=0$, $v=0$ at all the other boundaries. The streamfunction values are set to be zero at all the boundaries $(\psi=0)$. For this problem, the Reynolds number is defined as $Re=\dfrac{v_{0}R_{0}}{\nu}$, where $v_{0}$ is magnitude of the surface velocity of rotating side $DA$ and $\nu$ is the kinematic viscosity of the fluid. This is imperative to mention here that all boundary walls follow no-slip condition with the fluid inside the cavity, which supplies the following Neumann boundary conditions for vorticity.
@@
The computations for this problem are carried out for Reynolds numbers 55, 350, 1000, 2000, and 3000. The post processed steady-state streamfunction $(\psi)$ and vorticity $(\omega)$ contours for these $Re$ values are plotted in fig. 4. As $Re$ increases the primary vortex is seen to move gradually towards the right wall of the cavity. Similar to the case of flow inside a square cavity, with an increase in the value of $Re$ the secondary vortices at the corners opposite to the moving wall grow in size, particularly the secondary left vortex gains significantly more size as compared to the secondary right vortex and for higher values of $Re$ the former tends to occupy the entire left side of the cavity. Further, at $Re=3000$ a tertiary vortex is seen to form at the top left corner of the cavity. It is to be mentioned that all the findings are in accordance with those available in the literature [16, 9, 10, 17, 11] both experimentally and numerically. Additionally, at higher $Re$ values the vorticity contours move away from the center of the cavity toward the cavity walls thereby developing very strong vorticity gradients in the vicinity of the cavity walls.
@@ 55 $33^{2}$
Present -3.346e-3 0.720 1.741 -7.062e-3 -0.452 1.149
@@ 4.3 Problem 3: Natural convection in horizontal concentric annulus
We further analyze the efficiency of the newly developed scheme to tackle heat transfer by implementing it to solve natural heat convection inside a horizontal concentric annulus. This problem has gained a considerable amount of attention from researchers over the years due to its relevance and applicability in various engineering and physical situations, such as heat exchangers, solar collectors, nuclear reactors, thermal energy storage systems, etc [19, 20, 21, 22, 23, 24, 25, 26, 27]. It is governed by the transient Boussinesq equations which in the polar coordinate system are given as,
@@
The problem setup consists of two concentric circular walls of which the inner circle has a radius $R_{i}$ and the outer circle has a radius $R_{o}$. Two different temperatures are maintained at the inner and outer circular walls of the annulus as shown in fig. 10. The simulation is carried out for $Ra$ values $32.38e$ 3, $39.50e$ 3, $44.70e$ 4, $46.19e$ 4, and $51.02e$ 5. The values of the $Pr$ is set as 0.706, 0.717, and 0.718. For each case, we have considered $L/D_{i}$=0.8, where $L=R_{o}-R_{i}$ and $D_{i}=2\texttimes R_{i}$. The inner wall is heated to a unit nondimensional temperature $(\mathcal{T}_{h}=1)$ keeping the outer wall’s temperature at zero $(\mathcal{T}_{c}=0)$. This temperature difference serves as the driving force for the flow in this simulation. The values of streamfunction and vorticity at the inner and outer boundaries of the annulus are obtained respectively from no-slip criteria between the fluid and the walls of the annulus and one-sided approximations introduced in equations 31, 32, 33 and 34. The prior studies help us to identify that the near boundary region of the inner circle and the radial line $\theta=\dfrac{\pi}{2}$ should be observed cautiously. This motivates us to work with a mesh where grid points are clustered in these regions (see fig. 10). The grid is generated by making adequate changes in equations 25 and 35 and choosing the associated parameters as $L_{r}=0.8$, $L_{\theta}=2\pi$, $\Theta_{r}=\pi$, $\Theta_{\theta}=2\pi$, $\lambda_{r}=-0.4$ and $\lambda_{\theta}=0.25$.
@@
The numerical results of the present simulation have been presented in figs. 11 and 12. We carry out a qualitative comparison in fig. 11 where the numerically obtained steady-state isotherm contours for $Ra=4.7e4$, $Pr$=0.706 are depicted alongside that of the experimental study [21]. The figure shows that the fluid near the hot inner wall of the annulus flows up along the inner boundary and a strong plume emerges above the inner boundary at $\theta=\dfrac{\pi}{2}$, while the fluid near the cold outer wall flows down to the bottom of the annuli as an effect of buoyancy caused by the temperature difference. Once the flow reaches a steady state, the temperature field becomes symmetric. The maximum temperature gradient appears in the raised plume. Good agreements are obtained between the experimental and numerical solutions. Fig. 12 contains the steady-state isotherms alongside the streamfunctions for different $Ra$ values. The figure shows that for lower $Ra$ values natural convection is dominated by heat conduction. Almost concentric isotherms can be seen for $Ra=2.38e3$ while two symmetric vortices can be seen caused by the weak natural convection. As the value of $Ra$ increases the fluid motion driven by buoyancy force also increases, leading to a stronger convection. As a result, the isotherms start moving upward and change shape to form a plume. The plume becomes more prominent with a further increase in the $Ra$ value.